# Hydrodynamic noise in one dimension: projected Kubo formula and its vanishing in integrable models

### v1

@@ 3.2 Coarse graining and the origin of hydrodynamic noise
We now provide general arguments for why (3.1) holds. Although these general argument suggests the form of each correction, they do not allow us to compute them. However, Eq. (3.1), as a principle for the $-\ell$ asymptotic expansion of connected correlation functions, give us a solid framework from which we will evaluate à la Kubo, in the next section, all constitutive elements in terms of equilibrium correlation functions, in linearly degenerate systems.
Consider the fluid-cell mean of a conserved density $\overline{q_{i}}(x,t)$. By the conservation law, its time derivative, with respect to time $\mathrm{t}=\ell t$ at position ${\rm x}=\ell x$ in microscopic units, a is
$\frac{\partial}{\partial\mathrm{t}}\overline{q_{i}}(x,t)=L^{-1}(j_{i}({\rm x}-L/2,\ell t)-j_{i}({\rm x}+L/2,\ell t)).$ (27)
Assuming finite densities and finite typical microscopic velocities, the current observables are of order 1 and vary on a time scale $T$ of order 1 as well. Therefore, $\overline{q_{i}}(x,t)$ has variations of order $\mathcal{O}(L^{-1})$, and this happen on a time scale of order 1. That is, on $T=\mathcal{O}(1)$, $\overline{q_{i}}(x,t)$ can be considered to be invariant, and in particular non-fluctuating.
Now consider the fluid-cell mean of an observable that is not conserved, $\overline{o}(x,t)$. Because it is composed of $\mathcal{O}(L)$ essentially independent observables, each affected by independent fluctuations, the quantity $\overline{o}(x,t)$ varies on a time scale of order $1/L$, each variation being of order $1/L$. Therefore, $\overline{o}(x,t)$ fluctuates quickly, with $\mathcal{O}(L)$ variations, coming from “interaction events” such as collision of particles within the fluid cell, on a time $T$ of order 1. Averaging over $T$, which averages over these fluctuations, the fluid-cell mean of an arbitrary observable should relax to its microcanonical value $\mathsf{o}(\overline{q_{1}}(x,t),\overline{q_{2}}(x,t),\ldots)$.
@@
The above can only be true for $L$ large enough. The larger the mesoscopic length $L$, the more accurate is its microcanonical relaxation. Yet the cell size cannot be larger than the macroscopic variation scale $\ell$, because the relaxation argument made above requires spatial homogeneity. Further, the larger the time $T$, the more accurate is microcanonical relaxation as well, because more re-organisation has occurred. But $T$ must not bee too large so that conserved densities can be considered fixed and non-fluctuating. Our conjecture is that the optimal choice is obtained for
$LT=\mathcal{O}(\ell).$ (29)
For instance, $L=\ell^{1-\epsilon}$ and $T=\ell^{\epsilon}$ for $\epsilon>0$ small. The time $T$ is not finite, so the asymptotic expansion as $\ell\to\infty$ introduces mistakes of order $\epsilon$. The limit $\epsilon\to 0$ is then taken after the asymptotic expansion as $\ell\to\infty$. Another example is $L=\epsilon\ell$ and $T=1/\epsilon$. Here it is the fluid cell that is too large, introducing mistakes of order $\epsilon$, so that, again, the limit $\epsilon\to 0$ is taken after the asymptotic expansion in $\ell$. This is what is meant by the limit $\ell\gg L\gg\ell_{\rm micro}$.
What are the corrections at order $1/\ell$? The argument above suggests that there are three potential corrections.
Corrections to the microcanonical average. Because the cell size is not infinite, the exact average at fixed conserved densities is not the infinite-volume microcanonical average. Hence on the right-hand side of (28) there is a correction
$L^{-1}\delta\mathsf{o}(\overline{\underline{q}}(x,t))$ (30)
that is
@@
This is the finite-$L$ microcanonical average up to, including, $L^{-1}$ corrections. We will see in Sec. 4.2 that these corrections are indeed of order $L^{-1}$. The result appears to be non-universal, as, contrary to $\ell$ and $\ell_{\rm micro}$, the fluid cell size $L$ is not physical, but simply an artefact of our construction. However, the emergent fluctuation theory $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$ implicitly depends on $L$, and we will show in Sec. 4.2 how, in linearly degenerate systems, these $L$-dependences effectively cancel out and the result is, in fact, universal.
Diffusion. According to the argument above, there is a delay between the slow variations of the densities and relaxation, so that it is $\overline{o}(x,t+T/\ell)$ that takes the microcanonical value $\mathsf{o}^{\rm reg}(\overline{\underline{q}}(x,t))$. The $\ell^{-1}$ expansion gives an additional term, which is expected to be proportional to the spatial variations $\partial_{{\rm x}}\overline{q}_{i}(x,t)=\ell^{-1}\partial_{x}\overline{q}_{i}(x,t)$. Hence on the right-hand side of (28) there is another correction, for $t>0$:
$-\frac{1}{2}\ell^{-1}\sum_{i}\hat{\mathfrak{D}}_{o}^{~{}i}\big{(}\overline{\underline{q}}(x,t)\big{)}\,\partial_{x}\overline{q_{i}}(x,t)$ (32)
for some set of diffusion functions $\hat{\mathfrak{D}}_{o}^{~{}i}(\overline{\underline{q}}(x,t))$ associated to the observable $o$, where the factor $-\frac{1}{2}$ is conventional. For evolution towards $t<0$, we use PT symmetry, giving the factor $\operatorname{sgn}t$ as both $o$ and $q_{i}$’s are assumed to be PT invariant.
Emergent noise. Microcanonical relaxation, at finite $L$, receives stochastic corrections according to the central limit theorem. This is the result of forgetting the microscopic degrees of freedom when projecting onto the conserved densities. Hence on the right-hand side of (28) there is a noise correction,
@@
with the set of $\xi_{o}(x,t)$’s, for all local observables $o$’s, being correlated centered Gaussian white noises, with (20). The scaling in $\ell^{-1}$ is obtained as follows. The time-averaged fluid cell over $L\times T$ has $LT=\mathcal{O}(\ell)$ independent fluctuations, therefore a variance due to the noise of order $\mathcal{O}(\ell^{-1}$). Let us write a stochastic contribution $\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})$ in microscopic coordinates, representing the roughness in both space and time. This can be chosen centered, however it is not a white noise. Take its two-point correlations as $\langle\!\langle\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\xi_{o^{\prime}}^{\rm micro}(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\delta(\mathrm{x}-\mathrm{x}^{\prime})\delta(\mathrm{t}-\mathrm{t}^{\prime})$. Then, the variance of its total over a fluid cell is
@@
${\rm var}\Big{(}\frac{1}{LT}~\int_{-L/2}^{L/2}{\rm d}{\rm x}\int_{0}^{T}\,{\rm d}\mathrm{t}\,\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\Big{)}=\frac{1}{LT}\hat{\mathfrak{L}}_{o,o^{\prime}}=\ell^{-1}\hat{\mathfrak{L}}_{o,o^{\prime}}.$ (34)
@@
A similar calculation holds for covariances. Transforming to macroscopic coordinates, we obtain (33) with (20), and on that scale the noise becomes Gaussian as higher-point correlations are expected to vanish faster with $\ell^{-1}$.
4 Constitutive relations and the hydrodynamic equation
@@
We now evaluate the beyond-Euler components of the nonlinear fluctuating Boltzmann-Gibbs principle (3.1). For this purpose, we evaluate various equilibrium correlation functions. It turns out that in order to evaluate correlation functions, directly using (3.1), the asymptotic form of correlation functions in (22), and the cumulant expansion principles from Malyshev’s formula (see [26]), are sufficient. In particular, the Euler-scale projection formulas from [25, 26] for 2- and 3-point functions are sufficient for quantities in (3.1) with a factor of $\ell^{-1}$.
The results below hold in $d=1$ linearly degenerate systems [11, 12, 13] (see also [14]). These are systems which possess as many Riemann invariants $n_{I}$ as conserved quantities, that is, the Euler-scale hydrodynamic equation can be written in the form (see Appendix A)
$\partial_{t}n_{I}+\mathsf{v}_{I}\partial_{x}n_{I}=0,$ (35)
and such that the hydrodynamic velocity of mode $I$ does not depend on mode $I$ itself,
$\frac{\partial\mathsf{v}_{I}}{\partial n_{I}}=0\quad\forall\;I.$ (36)
In such systems, no shock appear [15, 16], hence solutions to the Euler equation are expected to be unique at all times. But most importantly for our purposes, in such systems:
@@
1. Correlation functions of conserved densities keep their local-equilibrium form even in non-equilibrium states such as those coming from the long-wavelength initial state (13),
$\langle\!\langle q_{i_{1}}(x_{1},t),q_{i_{2}}(x_{2},t)\rangle\!\rangle_{\ell}^{\rm c}=\mathsf{C}_{i_{1},i_{2}}(x_{1},t)\delta(x_{1}-x_{2})+\mbox{regular}.$ (37)
@@
Here the equilibrium covariance matrix is given in (10), and it is to be evaluated at the state characterised by $\underline{\mathsf{q}}(x_{1},t)=\langle\!\langle\underline{q}(x_{1},t)\rangle\!\rangle_{\ell}$. Eq. (37) was shown to hold in linearly degenerate systems using the BMFT in [25] and [19, Suppl Mat].
@@
2. For every local observable, the three-point couplings, Eq. (106), in space-time stationary, maximal entropy states vanish for modes of equal velocities:
@@
$\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}_{\underline{\beta}}=0\quad\mbox{if}\ \mathsf{v}_{I}=\mathsf{v}_{I^{\prime}}.$ (38)
@@
This implies that there is no superdiffusion [9, 10], and was shown from linear degeneracy in [19], Supplementary Material Sec. “Degenerate three-point coupling from linear degeneracy”.
Using these properties of linearly degenerate systems, we will show:
@@
    Point-splitting regularisation. The microcanonical regularisation is the point-splitting regularisation introduced in [19],
    $\Big{[}q_{i_{1}}(x_{1},t_{1})\cdots q_{i_{k}}(x_{k},t_{k})\Big{]}^{\rm reg}=\frac{1}{k!}\sum_{\sigma\in S(k)}q_{i_{1}}(x_{1}+\sigma_{i}0^{+},t_{1})\cdots q_{i_{k}}(x_{k}+\sigma_{k}0^{+},t_{k})$ (39)
    where $S(k)$ is the set of permutations of $(1,2,\ldots,k)$, and $0^{+}$ is a fixed, positive infinitesimal number.
@@
    Projected Kubo formula. The noise covariance coefficient is given by the projected Kubo formula: with $\langle\cdots\rangle=\langle\cdots\rangle_{\underline{\beta}(\underline{\mathsf{q}})}$ the stationary, maximal entropy, macrocanonical ensemble with averages $\langle q_{i}\rangle=\mathsf{q}_{i}$ of conserved densities,
@@
    Here $Q_{I}$ are the total conserved quantities in the normal mode basis (see (110)), $\mathsf{v}_{I}$ are the associated hydrodynamic velocities, and $\mathfrak{L}_{o_{1},o_{2}}$ is the conventional Onsager matrix
@@
    where the last equality follows by a general theorem [41].
@@
    Extended Einstein relation. The diffusion coefficients $\hat{\mathfrak{D}}_{o}^{~{}k}(\underline{\mathsf{q}})$ are related to the noise covariance by an extended Einstein relation
@@
It is also instructive to evaluate the Drude weight (42) within our formalism, in order to confirm that it is consistent. This is done in Appendix B.
@@
Our main result is Point II, the projected Kubo formula. Our proof uses a calculation that follows somewhat that done in [27]: the “diffusion from convection” term is subtracted. The subtraction corresponds to projecting out the quadratic charges forming the “scattering subspace” of the “diffusive space” constructed in [10], $\mathcal{H}_{\rm scat}\subset\mathcal{H}_{\rm dif}$. That is, with the projection $\mathbb{P}_{\rm scat}:\mathcal{H}_{\rm scat}\to\mathcal{H}_{\rm dif}$:

### v2

@@ 3.2 Coarse graining and the origin of hydrodynamic noise
Figure 1: A fluid cell of length $L$ over time $T$. There are any more interactions within the cell than there are boundary effects, leading to a separation of scales between fluctuations of arbitrary local observables and fluctuations of conserved densities.
We now provide general heuristic arguments for why (21) holds. Although these general arguments suggest the form of each correction, they do not allow us to compute them. However, Eq. (21), as a principle for the large-$\ell$ asymptotic expansion of connected correlation functions, gives us a framework from which we will evaluate à la Kubo, in the next section, all constitutive elements in terms of equilibrium correlation functions, in linearly degenerate systems.
Recall the fluid-cell mean (17). The main argument is that, because of the conservation laws, fluid-cell means of conserved densities are affected only by boundary effects during the dynamics, while generic observables are affected by every interactions within the volume of the cell. This leads to a separation of scales between conserved densities and generic observables. See Fig. 1.
Let us be more quantitative. By the conservation law, the time derivative of the fluid-cell mean of a conserve density, with respect to time $\mathrm{t}=\ell t$ at position ${\rm x}=\ell x$ in microscopic units, is
$\frac{\partial}{\partial\mathrm{t}}\overline{q_{i}}(x,t)=L^{-1}\big{(}j_{i}({\rm x}-L/2,\mathrm{t})-j_{i}({\rm x}+L/2,\mathrm{t})\big{)}.$ (33)
Assuming finite densities and finite typical microscopic velocities, the current observables are of order 1 and vary on a time scale $T$ of order 1 as well. Therefore, $\overline{q_{i}}(x,t)$ has variations of order $\mathcal{O}(L^{-1})$, and this happen on a time scale of order 1. Physically, these are due to surface effects, such as particles entering or exiting the fluid cell, and interaction between particles through the interface of the cell (which are the points ${\rm x}\pm L/2$); such surface effects are small. Then, on $T=\mathcal{O}(1)$, $\overline{q_{i}}(x,t)$ can be considered to be invariant, and in particular non-fluctuating.
Now consider the fluid-cell mean of an observable that is not conserved, $\overline{o}(x,t)$. Because it is composed of $\mathcal{O}(L)$ essentially independent observables, each affected by independent fluctuations, the quantity $\overline{o}(x,t)$ varies on a time scale of order $1/L$, each variation being of order $1/L$. Therefore, $\overline{o}(x,t)$ fluctuates quickly, with $\mathcal{O}(L)$ fluctuations on a time $T$ of order 1. Physcailly, these are due to volume effects: every “interaction event” between particles, such as collision of particles, within the volume of the cell, affect the fluid-cell mean of the observable, and on a time of order 1, there are $\mathcal{O}(L)$ such interaction events. They can be seen as independent fluctuations as they occur at random places within the volume of the cell and change the fluid-cell mean in incoherent ways. Averaging over $T$, which averages over these fluctuations, and considering that conserved densities are invariant and non-fluctuating, the fluid-cell mean of an arbitrary observable should therefore relax to its microcanonical value $\mathsf{o}(\overline{q_{1}}(x,t),\overline{q_{2}}(x,t),\ldots)$. See Fig. 2.
Figure 2: A fluid cell with the typical particle trajectories within it over a time $T$. There are $\mathcal{O}(LT)$ crossings, displayed as red stars, roughly representing independent interactions within this space-time region, while only $\mathcal{O}(T)$ boundary effects.
@@
The above can only be true for $\ell$ large enough. There are two concepts at play. The larger the mesoscopic length $L$ (here we follow the numbering below): (1) the closer is the average itself to the infinite-volume microcanonical average, which equals the macrocanonical average under the right identification of generalised temperatures (Sec. 2); and (3) the more fluctuations are taken into account, hence the less random corrections there are. Yet the cell size cannot be larger than the macroscopic variation scale $\ell$, because the relaxation argument made above requires spatial homogeneity.
Similarly, the larger the time $T$, the more accurate is microcanonical relaxation as well, because more re-organisation has occurred, and also (2) the less there is a time delay between the effective time at which relaxation has occured, and the time at which the values of the conserved densities are taken to determine the state towards which relaxation occurs. But $T$ must not bee too large so that conserved densities can be considered fixed and non-fluctuating.
Our conjecture is that the optimal choice is obtained for
$LT/\ell_{\rm micro}=\mathcal{O}(\ell).$ (35)
Note that neither $L$ nor $T$ are univeral quantity – they are artefact of our fluid-cell construction. But $\ell$ is physical – it is the physical variation length scale. Below we often simply set $\ell_{\rm micro}=1$, which we can do without loss of generality – this sets the scale of our microscopic coordinates.
For instance, $L=\ell^{1-\epsilon}$ and $T=\ell^{\epsilon}$ for $\epsilon>0$ small. The time $T$ is not finite but very small on the scale $\ell$, so the asymptotic expansion as $\ell\to\infty$ introduces small mistakes of order $\epsilon$. The limit $\epsilon\to 0$ is then taken after the asymptotic expansion as $\ell\to\infty$. Another example is $L=\epsilon\ell$ and $T=1/\epsilon$. Here it is the fluid cell that is too large, introducing mistakes of order $\epsilon$, so that, again, the limit $\epsilon\to 0$ is taken after the asymptotic expansion in $\ell$. This is what is meant by the limit $\ell\gg L\gg\ell_{\rm micro}$.
What are the corrections at order $1/\ell$ or $1/L$ beyond the $\ell\to\infty$ limit? The argument above suggests that there are three potential corrections.
1. Corrections to the microcanonical average, $L<\infty$. Because the cell size is not infinite, the exact average at fixed conserved densities is not the infinite-volume microcanonical average. Hence on the right-hand side of (34) there is a correction
$L^{-1}\delta\mathsf{o}(\overline{\underline{q}}(x,t)).$ (36)
We gather this into
@@
This is the finite-$L$ microcanonical average up to, including, $L^{-1}$ corrections. We will see in Sec. 4.1 that these corrections are indeed of order $L^{-1}$. The result appears to be non-universal; as mentioned, contrary to $\ell$, the fluid cell size $L$ is not physical but simply an artefact of our construction. However, the emergent fluctuation theory $\langle\!\langle\cdots\rangle\!\rangle_{\ell}$ implicitly depends on $L$ because it is a theory for fluctuations of fluid-cell means, and we will show in Sec. 4.1 how, in linearly degenerate systems, these $L$-dependences effectively cancel out and the result is, in fact, universal.
2. Time delay, $T>0$: diffusion term. According to the argument above, there is a delay between the slow variations of the densities and relaxation, so that it is $\overline{o}(x,t+T/\ell)$ that takes the microcanonical value $\mathsf{o}^{\rm reg}(\overline{\underline{q}}(x,t))$. The $\ell^{-1}$ expansion gives an additional term. This is expected to be proportional to the spatial variations $\partial_{{\rm x}}\overline{q}_{i}(x,t)=\ell^{-1}\partial_{x}\overline{q}_{i}(x,t)$, because $L$ must be taken smaller if the variations are larger, thus $T$ must be taken larger to keep $LT$ the same. Hence on the right-hand side of (34) there is another correction, for $t>0$:
$\frac{1}{2}\ell^{-1}\sum_{i}\hat{\mathfrak{D}}_{o}^{~i}\big{(}\overline{\underline{q}}(x,t)\big{)}\,\partial_{x}\overline{q_{i}}(x,t)$ (38)
for some set of diffusion functions $\hat{\mathfrak{D}}_{o}^{~i}(\overline{\underline{q}}(x,t))$ associated to the observable $o$, where the factor $\frac{1}{2}$ is conventional. For evolution backwards in time, $t<0$, we use PT symmetry, giving the factor $\operatorname{sgn}t$ as both $o$ and $q_{i}$’s are assumed to be PT invariant. Note how the direction of time is fundamental in this part of the argument: there is a delay in reaching the microcanonical average.
3. Emergent noise, $LT<\infty$. Microcanonical relaxation, at finite $L$, receives stochastic corrections according to the central limit theorem. Intuitively, this is the result of partially remembering the microscopic degrees of freedom that have been projected out when projecting onto the conserved densities in the nonlinear Boltzmann-Gibbs principle. Hence on the right-hand side of (34) there is a noise correction,
@@
with the set of $\xi_{o}(x,t)$’s, for all local observables $o$’s, being centered Gaussian white noises, delta-correlation in space-time. We need to explain three aspects: (a) why the scaling is $\ell^{-1}$; (b) why the emergent noise is delta-correlated in space-time; and (c) why it is the Itô convention that must be used.
@@
(a) The scaling in $\ell^{-1}$ is obtained as follows. The time-averaged fluid cell over $L\times T$ has $LT=\mathcal{O}(\ell)$ independent fluctuations, therefore a variance due to the noise of order $\mathcal{O}(\ell^{-1}$). Let us write a stochastic contribution $\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})$ in microscopic coordinates, representing the roughness in both space and time. This can be chosen centered, however it is not Gaussian because it encodes small-scale effects, and it is not a white noise because there are correlations on lengths $\mathcal{O}(\ell_{\rm micro})=\mathcal{O}(1)$. We now argue that its two-point correlations can be taken as $\langle\!\langle\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\xi_{o^{\prime}}^{\rm micro}(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\delta_{\ell_{\rm micro}}(\mathrm{x}-\mathrm{x}^{\prime},\mathrm{t}-\mathrm{t}^{\prime})$ for some $\hat{\mathfrak{L}}_{o,o^{\prime}}$, with mollified space-time delta-function on scale the linear $\ell_{\rm micro}$. Indeed, with this, the variance of its mean over a fluid cell in space-time is then $\mathcal{O}(\ell^{-1})$, as it should:
@@
${\rm var}\Big{(}\frac{1}{LT}\nobreakspace\int_{-L/2}^{L/2}{\rm d}{\rm x}\int_{0}^{T}\,{\rm d}\mathrm{t}\,\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\Big{)}\sim\frac{1}{LT}\hat{\mathfrak{L}}_{o,o^{\prime}}=\mathcal{O}(\ell^{-1})\qquad(\ell\to\infty).$ (40)
@@
A similar calculation holds for covariances. Transforming to macroscopic coordinates $\delta_{\ell_{\rm micro}}(\ell x-\ell x^{\prime},\ell t-\ell t^{\prime})\to\ell^{-2}\delta(x-x^{\prime})\delta(t-t^{\prime})$, we write $\xi_{o}^{\rm micro}(\ell x,\ell t)=\ell^{-1}\xi_{o}(x,t)$ and we obtain (39) with (32), and on that scale the noise becomes Gaussian as higher-point correlations are expected to vanish faster with $\ell^{-1}$.
(b) The delta-correlation in space-time is argued for as follows. It is known that correlations of physical observables in space (out of equilibrium) and in time (in and out of equilibrium) can be strong, and that this is due to hydrodynamic mode propagation. Since hydrodynamic modes are waves of conserved densities, such strong correlations are destroyed when fixing the space-time configuration of conserved densities on every fluid cell. There remains only short-range correlations on scale $\ell_{\rm micro}$, thus giving $\langle\!\langle\xi_{o}^{\rm micro}(\mathrm{x},\mathrm{t})\xi_{o^{\prime}}^{\rm micro}(\mathrm{x}^{\prime},\mathrm{t}^{\prime})\rangle\!\rangle_{\ell}=\hat{\mathfrak{L}}_{o,o^{\prime}}\delta_{\ell_{\rm micro}}(\mathrm{x}-\mathrm{x}^{\prime},\mathrm{t}-\mathrm{t}^{\prime})$.
(c) Finally, the Itô convention is again an effect of time delay. Fluctuations occur on a time following that at which the essentially fixed fluid-cell means of conserved densities are taken. Thus (28) holds, where Gaussian forward increments have a variance that is determined by the conserved densities at the initial time of the increment. Likewise (32) holds thanks to this.
Remark 3.3.
In stochastic systems, the emergent noise can often be obtained in a mathematically accurate fashion. But such an understanding is still lacking in general in Hamiltonian systems.
4 Constitutive relations
@@
We now evaluate the beyond-Euler components of the nonlinear fluctuating Boltzmann-Gibbs principle (21). For this purpose, we evaluate various equilibrium correlation functions. It turns out that in order to evaluate correlation functions directly using (21), it is sufficient to know the asymptotic form (20), and to use the cumulant expansion principles from Malyshev’s formula (see [43]). In particular, the Euler-scale projection formulas from [8, 43] for 2- and 3-point functions are sufficient for quantities in (21) with a factor of $\ell^{-1}$.
The results below hold in $d=1$ “no-shock” systems which are strictly hyperbolic. Strict hyperbolicity means that all hydrodynamic velocities are distinct. Further, a no-shock system is a many-body system whose Euler-scale equation does not develop shock in non-equilibrium situations. In such systems the solutions to Euler equations are unique at all times, and nonlinear Euler-scale projections are valid everywhere in space-time [43]. The most important examples (perhaps the only ones!) of such systems are linearly degenerate systems [26, 27, 28, 29, 30], as indeed these do not develop shocks [31, 32, 30]. See Appendix C for the definition of such systems. It is also known that no shocks are produced in integrable systems [34, 35], which are, in a formal sense, also linearly degenerate [33].
We show in Appendix C that, in no-shock systems, the following results hold:
@@
1. Correlation functions of conserved densities keep their local-equilibrium form at later time even in long-wavelength non-equilibrium states such as those emanating from (12), at the Euler scale:
$\langle\!\langle q_{i_{1}}(x_{1},t),q_{i_{2}}(x_{2},t)\rangle\!\rangle_{\ell}^{\rm c}=\ell^{-1}\Big{(}\mathsf{C}_{i_{1},i_{2}}(x_{1},t)\delta(x_{1}-x_{2})+\mbox{regular}\Big{)}+\mathcal{O}(\ell^{-2}).$ (41)
@@
Here the equilibrium covariance matrix is given in (9), and it is to be evaluated at the state characterised by $\underline{\mathsf{q}}(x_{1},t)=\langle\!\langle\underline{q}(x_{1},t)\rangle\!\rangle_{\ell}$. This is consistent with the initial condition (24). A partial proof for Eq. (41) was given using the BMFT in [8], and a related calculation in [25, Suppl Mat]. We provide a complete proof in Appendix C, which requires strict hyperbolicity.
@@
2. For every local observable, the three-point couplings, Eq. (137), in space-time stationary, maximal entropy states vanish for normal modes of equal velocities:
@@
$\langle Q_{I},Q_{I^{\prime}},o_{1}^{-}\rangle^{\rm c}_{\underline{\beta}}=0\quad\mbox{if}\ \mathsf{v}_{I}=\mathsf{v}_{I^{\prime}}.$ (42)
@@
Thus, as we assume strict hyperbolicity, vanishing occurs whenever $I\neq I^{\prime}$ (but the more general result shown here holds also without strict hyperbolicity). This implies that there is no superdiffusion [23, 24], and was shown from linear degeneracy in [25], Supplementary Material Sec. “Degenerate three-point coupling from linear degeneracy”. For readability, we reproduce this proof in Appendix C.
Using these properties of no-shock systems, we will show:
@@
    Point-splitting regularisation. The microcanonical regularisation $\mathsf{o}^{\rm reg}(\underline{q}(x,t))$ is given by the point-splitting prescription introduced in [25],
    $\Big{[}q_{i_{1}}(x,t)\cdots q_{i_{k}}(x,t)\Big{]}^{\rm reg}=\frac{1}{k!}\sum_{\sigma\in S(k)}q_{i_{1}}(x+\sigma_{i}0^{+},t)\cdots q_{i_{k}}(x+\sigma_{k}0^{+},t)$ (43)
    where $S(k)$ is the set of permutations of $(1,2,\ldots,k)$, and $0^{+}$ is a fixed, positive infinitesimal number. Note how this is a sum over various orderings in space, with infinitesimal displacements.
@@
    Projected Kubo formula. The noise covariance coefficient is given by the projected Kubo formula: with $\langle\cdots\rangle=\langle\cdots\rangle_{\underline{\mathsf{q}}}$ the space-time stationary maximal-entropy state with averages $\langle q_{i}\rangle=\mathsf{q}_{i}$,
@@
    Here $Q_{I}$ are the total conserved quantities in the normal mode basis (see (142)), $\mathsf{v}_{I}$ are the associated hydrodynamic velocities, and $\mathfrak{L}_{o_{1},o_{2}}$ is given by the conventional Green-Kubo formula
@@
    where the last equality follows from a general theorem [51]. Note that the quantity $\mathfrak{L}_{ik}:=\mathfrak{L}_{j_{i},j_{k}}$ is the usual Onsager matrix.
@@
    (Extended) Einstein relation. The diffusion coefficients $\hat{\mathfrak{D}}_{o}^{~k}(\underline{\mathsf{q}})$ are related to the noise covariance by an Einstein relation, here extended to diffusion for arbitrary observables $o$ instead of only currents,
@@
It is also instructive to evaluate the Drude weight (46) within our formalism, in order to confirm that it is consistent. This is done in Appendix D.
@@
Our main result is Point II, the projected Kubo formula. Our proof uses a calculation that follows somewhat that done in [44]: the “diffusion from convection” term is subtracted. Crucially, the subtraction corresponds to projecting out the quadratic charges forming the “scattering subspace” of the “diffusive space” constructed in [24], $\mathcal{H}_{\rm scat}\subset\mathcal{H}_{\rm dif}$. The diffusive space is a Hilbert space constructed with respect to the sesquilinear form given by the Green-Kubo formula
$(o_{1},o_{2})_{\rm dif}:=\mathfrak{L}_{o_{1},o_{2}}(\underline{\mathsf{q}})=\int_{-\infty}^{\infty}{\rm d}\mathrm{t}\,\Big{(}\int{\rm d}{\rm x}\,\langle o_{1}({\rm x},\mathrm{t}),o_{2}(0,0)\rangle^{\rm c}-\mathsf{D}_{o_{1},o_{2}}\Big{)}$ (48)
(recall that our observables $o_{1},o_{2}$ are assumed to be real or Hermitian, however this is easily extended by sesquilinearity). That is, the space $\mathcal{H}_{\rm dif}$ is obtained by a Gelfand-Naimark-Segal construction: it is easy to show that $(o_{1},o_{2})_{\rm dif}$ is positive semi-definite, so its null space is moded out and the resulting space of equivalence classes is completed with respect to the norm $||\cdot||_{\rm dif}=\sqrt{(\cdot,\cdot)_{\rm dif}}$. The scattering subspace is that spanned by particular quadratic elements parametrised by two normal mode indices $I,J$:
$\mathcal{H}_{\rm scat}:={\rm span}\Big{\{}w_{IJ}=\lim_{X\to\infty}\frac{|\mathsf{v}_{I}-\mathsf{v}_{J}|}{2X}\int_{-X}^{X}{\rm d}{\rm x}\,q_{I}({\rm x})q_{J}\Big{\}}\subset\mathcal{H}_{\rm dif}.$ (49)
The fact that these lie within $\mathcal{H}_{\rm dif}$ is nontrivial and argued for in [24]. With the projection $\mathbb{P}_{\rm scat}:\mathcal{H}_{\rm scat}\to\mathcal{H}_{\rm dif}$, it turns out that the second term in (44) is exactly [24]
$(\mathbb{P}_{\rm scat}o_{1},\mathbb{P}_{\rm scat}o_{2})=\frac{1}{2}\sum_{I,I^{\prime}:\mathsf{v}_{I}\neq\mathsf{v}_{I}^{\prime}}\frac{\langle o_{1}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}\langle o_{2}^{-},Q_{I},Q_{I^{\prime}}\rangle^{\rm c}}{|\mathsf{v}_{I}-\mathsf{v}_{I^{\prime}}|}.$ (50)
That is,